# Quantum Corrections to Symmetron Fifth-Force Profiles

### v1

@@ 3.1 Thin-Wall Approximation
We want an analytical understanding of the quantum corrections, but the above system of equations cannot be solved in closed form. One could obtain an approximate solution by linearising the equation of motion [51], but then the Green’s function would be independent of the field. Instead, we consider a source whose radius $R$ is large compared to the characteristic scale of variations of the classical field. That is, $\mu R\gg 1$. Then, $\phi$ remains small and grows slowly in the interval $0<r<R$, experiences a rapid jump around $r=R$ and then quickly flattens out to $v$ as $r\rightarrow\infty$. Thus, we may drop the damping term in eq. (3.10) and turn what was once a radial problem into a purely one-dimensional problem. It is perhaps more conceptually clear to think of this approximation as an asymptotic approximation in the limit $R\rightarrow\infty$. For notational clarity, we write s=rRs=r-R so that the equation of motion reads
@@
We denote by $\phi_{-}(s)$ and $\phi_{+}(s)$ the interior and exterior fields respectively. Separately solving for them is simply a matter of picking appropriate values for $s_{1}$ and $s_{2}$. The full profile,
$\phi(s)=\Theta(-s)\phi_{-}(s)+\Theta(s)\phi_{+}(s)\>,$ (3.17)
@@
for some real $c_{+}$. For the interior contribution, choosing $s_{1}=$-$\infty$ and $s_{2}=s<0$ implies
@@
$g=\sqrt{2\left(\frac{\rho_{0}}{\mu^{2}M^{2}}-1\right)}\>,$ (3.25)
is the effective mass of interior scalar fluctuations, in units of $\gamma$. By the identity,
@@
with $c_{-}\in\mathbb{R}$. Continuity with $\chi_{+}$ is easily achieved by defining the constants $c_{\pm}$ in terms of the value of the field at the boundary of the source $\chi_{0}$,
@@
In conclusion, the static classical field profile is exactly
$\varphi_{\text{cl}}(s)=v\left\{\Theta(-s)g\,\text{csch}\>\!\left[\text{arcsch}\>\!\left(\frac{\chi_{0}}{g}\right)-g\gamma s\right]+\Theta(s)\tanh\left(\gamma s+\text{artanh}\>\!\chi_{0}\right)\right\}\>,$ (3.30)
@@ 4 Green’s Function
We now turn to the calculation of the Green’s function. Through the use of the thin-wall approximation, we again reduce the spherical problem to a one-dimensional one. As one would expect from the problem of eigenmodes of a system with a piecewise continuous mass (see, for example, Appendix A), the linearly independent solutions which compose the Green’s function behave differently for different frequency intervals. One can regard these expressions as equivalent following a Feynman contour prescription.
4.1 Frequency Domain
The differential operator from which the Green’s function derives is related to the second functional derivative evaluated on $\varphi_{\text{cl}}$, slightly altered from the more general (2.10),
@@ 4.1 Frequency Domain
We refer to the operator in parentheses as the fluctuation operator,
@@
where $P_{l}$ is the Legendre polynomial of order $l$, allowing us to write
@@
Substituting this into eq. (4.7), and taking the thin-wall approximation (which amounts to dropping the damping term and replacing $r\rightarrow R$ in the centrifugal potential and discontinuity [40]), yields
@@
$G^{>}_{l}(s,s^{\prime};E)=G^{<}_{l}(s^{\prime},s;E)\>,$ (4.11)
it follows that $G_{l}(s,s^{\prime};E)$, which is given by
$G_{l}(s,s^{\prime};E)=\Theta(s^{\prime}-s)G^{<}_{l}(s,s^{\prime};E)+\Theta(s-s^{\prime})G^{>}_{l}(s,s^{\prime};E)\>,$ (4.12)
$G^{\gtrless}_{l}(s,s^{\prime};E)=\Theta_{s}\Theta_{s^{\prime}}G_{l,+}^{\gtrless}(s,s^{\prime};E)+\Theta_{s}\Theta_{-s^{\prime}}G_{l,\pm}^{\gtrless}(s,s^{\prime};E)+\Theta_{-s}\Theta_{-s^{\prime}}G_{l,-}^{\gtrless}(s,s^{\prime};E)\>,$ (4.13)
where we append a sign to the subscript to identify these contributions. In order of appearance, we refer to each term as the exterior, boundary and interior contributions. A graphical summary of this decomposition is given in Figure 3. We will solve for contributions to the $G_{l}^{\gtrless}$ for different values of $E$, adopting a naming scheme analogous to that in quantum mechanics: contributions for $0<E<2\gamma$ are called bound, $2\gamma<E<g\gamma$ tunnelling and $E>g\gamma$ scattering.
[…]
Bound Regime
When $0<E<2\gamma$, the differential equation
$\left[-\frac{\text{d}^{2}}{\text{d}s^{2}}+\frac{l(l+1)}{R^{2}}-E^{2}-2\gamma^{2}+6\gamma^{2}\chi_{+}(s)^{2}\right]y_{+}(s)=0$ (4.14)
$u=\chi_{+}=\tanh\left(\gamma s+c_{1}\right)\Rightarrow\frac{\text{d}}{\text{d}s}=\gamma\left(1-\chi_{+}^{2}\right)\frac{\text{d}}{\text{d}\chi_{+}}\>,$ (4.15)
we transform this equation to
$\frac{\text{d}}{\text{d}u}\left[\left(1-u^{2}\right)\frac{\text{d}y_{+}}{\text{d}u}\right]+\left[j(j+1)-\frac{n^{2}}{1-u^{2}}\right]y_{+}(u)=0\>,$ (4.16)
and recover the general Legendre equation with degree $j=2$ and order
$n=\frac{1}{\gamma}\sqrt{\frac{l(l+1)}{R^{2}}-E^{2}+4\gamma^{2}}\>,$ (4.17)
$P_{j}^{n}(x)=\sec(n\pi)\frac{\Gamma(j+n+1)}{\Gamma(j-n+1)}P_{j}^{-n}(x)+\frac{2}{\pi}\tan(n\pi)Q_{j}^{n}(x)\>,$ (4.18)
so we may write
$y_{+}(s)\sim P_{2}^{\pm n}\left(u(s)\right)\>.$ (4.19)
The equation
$\left[-\frac{\text{d}^{2}}{\text{d}s^{2}}+\frac{l(l+1)}{R^{2}}-E^{2}+g^{2}\gamma^{2}+6\gamma^{2}\chi_{-}(s)^{2}\right]y_{-}(s)=0$ (4.20)
$\upsilon=\frac{\chi_{-}}{g}=\text{csch}\>\!(c_{2}-g\gamma s)\Rightarrow\frac{\text{d}}{\text{d}s}=g\gamma\upsilon\sqrt{1+u^{2}}\frac{\text{d}}{\text{d}\upsilon}\>,$ (4.21)
transforming the equation into
$\left[\upsilon^{2}(1+\upsilon^{2})\frac{\text{d}^{2}}{\text{d}\upsilon^{2}}+\upsilon(1+2\upsilon^{2})\frac{\text{d}}{\text{d}\upsilon}+\frac{4-n^{2}-g^{2}}{g^{2}}-6u^{2}\right]y_{-}(\upsilon)=0\>.$ (4.22)
This equation bears a strong similarity to the hypergeometric equation in $\upsilon^{2}$. To see this more clearly, we use the change of variables $z=\upsilon^{2}$ to write
$\left[z(1+z)\frac{\text{d}^{2}}{\text{d}z^{2}}+\left(1+\frac{3}{2}z\right)\frac{\text{d}}{\text{d}z}-\frac{3}{2}+\frac{4-n^{2}-g^{2}}{4g^{2}z}\right]y_{-}(z)=0\>,$ (4.23)
then “peel off” the large $z$ behaviour with the substitution
$y_{-}(z)=z^{a}f(z;a)\>,$ (4.24)
where $a$ is a free parameter and $f$ may depend on $a$. Choosing $a$ to be
$a=\frac{1}{2g\gamma}\sqrt{\frac{l(l+1)}{R^{2}}-E^{2}+g^{2}\gamma^{2}}\>,$ (4.25)
allows us to write the solutions in terms of hypergeometric functions
$f_{1}(z,a)={}_{2}F_{1}\left(a-1,a+\frac{3}{2};2a+1;-z\right)$ (4.26)
and
$f_{2}(z;a)=z^{-2a}{}_{2}F_{1}\left(-a-1,-a+\frac{3}{2};-2a+1;-z\right)\>.$ (4.27)
$F_{a}(z)=z^{a}{}_{2}F_{1}\left(a-1,a+\frac{3}{2};2a+1;-z\right)\>,$ (4.28)
and write
$y_{-}(s)\sim F_{\pm a}\left(z(s)\right)\>.$ (4.29)
Note that $a$ will remain real until we reach the scattering regime.
Let $u^{\prime}=u(s^{\prime})$ and $z^{\prime}=z(s^{\prime})$:
Exterior:
$G_{l,+}^{>}(s,s^{\prime};E)=P_{2}^{-n}(u)\left(AP_{2}^{n}(u^{\prime})+BP_{2}^{-n}(u^{\prime})\right)=G_{l,+}^{<}(s^{\prime},s;E)\>,$ (4.30)
where $A$ and $B$ are constants.
Boundary:
$G^{>}_{l,\pm}(s,s^{\prime};E)=\frac{P_{2}^{-n}(u)F_{a}(z^{\prime})}{\mathcal{W}}=G^{<}_{l,\pm}(s^{\prime},s;E)\>,$ (4.31)
where $\mathcal{W}$ is a constant.
Interior:
$G_{l,-}^{>}(s,s^{\prime};E)=F_{a}(z^{\prime})\left(CF_{-a}(z)+DF_{a}(z)\right)=G_{l,-}^{<}(s^{\prime},s;E)\>,$ (4.32)
where $C$ and $D$ are both constants.
$\left.\frac{\text{d}}{\text{d}s}G^{>}(s,s^{\prime})\right|_{s=s^{\prime}}-\left.\frac{\text{d}}{\text{d}s}G^{<}(s,s^{\prime})\right|_{s=s^{\prime}}=\frac{1}{R^{2}}\>.$ (4.33)
$A=\frac{1}{R^{2}}\left[W\left(P_{2}^{n}(u),P_{2}^{-n}(u)\right)\right]^{-1}=-\frac{\pi}{2\gamma R^{2}}\csc\left(n\pi\right)\>,$ (4.34)
$C=\frac{1}{R^{2}}\left[W\left(F_{a}(z),F_{-a}(z)\right)\right]^{-1}=\frac{1}{4ag\gamma R^{2}}\>,$ (4.35)
and
$\mathcal{W}=W(F_{a}(z_{0}),P_{2}^{-n}(u_{0}))R^{2}\>.$ (4.36)
$\frac{1}{\mathcal{W}R^{2}}=F_{a}(z_{0})^{-1}\left(-\frac{\pi}{2\gamma R^{2}}\csc\left(n\pi\right)P_{2}^{n}(u_{0})+BP_{2}^{-n}(u_{0})\right)$ (4.37)
while the second one implies
$\frac{1}{\mathcal{W}R^{2}}=\left.\left(\frac{\text{d}}{\text{d}s^{\prime}}F_{a}(z^{\prime})\right)^{-1}\left(-\frac{\pi}{2\gamma R^{2}}\csc\left(n\pi\right)\frac{\text{d}}{\text{d}s^{\prime}}P_{2}^{n}(u^{\prime})+B\frac{\text{d}}{\text{d}s^{\prime}}P_{2}^{-n}(u^{\prime})\right)\right|_{s^{\prime}=0}\>.$ (4.38)
$B=-\frac{\pi}{2\gamma R^{2}}\csc(n\pi)\frac{W\left(P_{2}^{n}(u_{0}),F_{a}(z_{0})\right)}{W\left(F_{a}(z_{0}),P_{2}^{-n}(u_{0})\right)}\>.$ (4.39)
$\frac{1}{\mathcal{W}R^{2}}=\left(P_{2}^{-n}(u_{0})\right)^{-1}\left(-\frac{1}{4ag\gamma R^{2}}F_{-a}(z_{0})+DF_{a}(z_{0})\right)\>,$ (4.40)
while the second one implies
$\frac{1}{\mathcal{W}R^{2}}=\left.\left(\frac{\text{d}}{\text{d}s}P_{2}^{-n}(u)\right)^{-1}\left(-\frac{1}{4ag\gamma R^{2}}\frac{\text{d}}{\text{d}s}F_{-a}(z)+D\frac{\text{d}}{\text{d}s}F_{a}(z)\right)\right|_{s=0}\>.$ (4.41)
$D=-\frac{1}{4ag\gamma R^{2}}\frac{W\left(P_{2}^{-n}(u_{0}),F_{-a}(z_{0})\right)}{W\left(F_{a}(z_{0}),P_{2}^{-n}(u_{0})\right)}\>.$ (4.42)
Let’s summarise our results. The exterior contributions are
$G_{l,+}^{>}(s,s^{\prime};E)=-\frac{\pi}{2\gamma R^{2}}\csc\left(n\pi\right)P_{2}^{-n}(u)\left(P_{2}^{n}(u^{\prime})+\frac{W\left(P_{2}^{n},F_{a}\right)}{W\left(F_{a},P_{2}^{-n}\right)}P_{2}^{-n}(u^{\prime})\right)=G_{l,+}^{<}(s^{\prime},s;E)\>,$ (4.43)
the interior contributions are
$G_{l,-}^{>}(s,s^{\prime};E)=-\frac{1}{4ag\gamma R^{2}}F_{a}(z^{\prime})\left(F_{-a}(z)+\frac{W\left(P_{2}^{-n},F_{-a}\right)}{W\left(F_{a},P_{2}^{-n}\right)}F_{a}(z)\right)=G_{l,-}^{<}(s,s^{\prime};E)\>,$ (4.44)
and the boundary contributions are
$G^{>}_{l,\pm}(s,s^{\prime};E)=\frac{P_{2}^{-n}(u)F_{a}(z^{\prime})}{W(F_{a},P_{2}^{-n})R^{2}}=G^{<}_{l,\pm}(s^{\prime},s;E)\>.$ (4.45)
All Wrońskians are evaluated at $s=0$.
Tunnelling Regime
$y_{+}(s)\sim P_{2}^{\pm i\nu}(u(s))\>,$ (4.46)
where
$\nu=\frac{1}{\gamma}\sqrt{E^{2}-\frac{l(l+1)}{R^{2}}-4\gamma^{2}}\>.$ (4.47)
Exterior:
$G^{>}_{l,+}(s,s^{\prime};E)=AP_{2}^{i\nu}(u)P_{2}^{-i\nu}(u^{\prime})+BP_{2}^{i\nu}(u)P_{2}^{i\nu}(u^{\prime})=G^{<}_{l,+}(s^{\prime},s;E)\>.$ (4.48)
Boundary:
$G^{>}_{l,\pm}(s,s^{\prime};E)=\frac{P_{2}^{i\nu}(u)F_{a}(z^{\prime})}{\mathcal{W}}=G^{<}_{l,\pm}(s^{\prime},s;E)\>.$ (4.49)
Interior:
$G_{l,-}^{>}(s,s^{\prime};E)=F_{a}(z^{\prime})\left(CF_{-a}(z)+DF_{a}(z)\right)=G_{l,-}^{<}(s^{\prime},s;E)\>.$ (4.50)
The derivative jump condition [eq. (4.33)] gives us three of the five constants,
$A=-\frac{i\pi}{2\gamma R^{2}}\operatorname{csch}(\nu\pi)\>,\,\mathcal{W}=W\left[F_{a}(z_{0}),P_{2}^{i\nu}(u_{0})\right]\>,\,C=\frac{1}{4ag\gamma R^{2}}\>,$ (4.51)
while continuity and differentiability give us the remaining two:
$B=-\frac{i\pi}{2\gamma R^{2}}\operatorname{csch}(\nu\pi)\frac{W\left(P_{2}^{-i\nu}(u_{0}),F_{a}(z_{0})\right)}{W\left(F_{a}(z_{0}),P_{2}^{i\nu}(u_{0})\right)}\>,\,D=\frac{1}{4ag\gamma R^{2}}\frac{W\left(P_{2}^{i\nu}(u_{0}),F_{-a}(z_{0})\right)}{W\left(F_{a}(z_{0}),P_{2}^{i\nu}(u_{0})\right)}\>.$ (4.52)
$G^{>}_{l,+}(s,s^{\prime};E)=-\frac{i\pi}{2\gamma R^{2}}\operatorname{csch}(\nu\pi)P_{2}^{i\nu}(u)\left(P_{2}^{-i\nu}(u^{\prime})+\frac{W\left(P_{2}^{-i\nu},F_{a}\right)}{W\left(F_{a},P_{2}^{i\nu}\right)}P_{2}^{i\nu}(u^{\prime})\right)\>,$ (4.53)
$G^{>}_{l,\pm}(s,s^{\prime};E)=\frac{P_{2}^{i\nu}(u)F_{a}(z^{\prime})}{W\left(F_{a},P_{2}^{i\nu}\right)R^{2}}\>,$ (4.54)
and
$G_{l,-}^{>}(s,s^{\prime};E)=\frac{1}{4ag\gamma R^{2}}F_{a}(z^{\prime})\left(F_{-a}(z)+\frac{W\left(P_{2}^{i\nu},F_{-a}\right)}{W\left(F_{a},P_{2}^{i\nu}\right)}F_{a}(z)\right)\>.$ (4.55)
Scattering Regime
$y_{-}(s)\sim F_{\pm i\alpha}(z(s))\>,$ (4.56)
where
$\alpha=\frac{1}{2g\gamma}\sqrt{E^{2}-\frac{l(l+1)}{R^{2}}-g^{2}\gamma^{2}}\>.$ (4.57)
Exterior:
The form of the exterior solutions does not change.
$G^{>}_{l,+}(s,s^{\prime};E)=AP_{2}^{i\nu}(u)P_{2}^{-i\nu}(u^{\prime})+BP_{2}^{i\nu}(u^{\prime})P_{2}^{i\nu}(u)=G^{<}_{l,+}(s^{\prime},s;E)\>.$ (4.58)
Boundary:
The boundary contributions are
$G^{>}_{l,\pm}(s,s^{\prime};E)=\frac{P_{2}^{i\nu}(u)F_{-i\alpha}(z^{\prime})}{\mathcal{W}}=G^{<}_{l,\pm}(s^{\prime},s;E)\>.$ (4.59)
Interior:
The interior contributions are
$G^{>}_{l,-}(s,s^{\prime};E)=CF_{-i\alpha}(z^{\prime})F_{i\alpha}(z)+DF_{-i\alpha}(z^{\prime})F_{-i\alpha}(z)=G^{<}_{l,-}(s^{\prime},s;E)\>.$ (4.60)
$A=-\frac{i\pi}{2\gamma R^{2}}\operatorname{csch}(\nu\pi)\>,\,\mathcal{W}=W\left(F_{-i\alpha}(z_{0}),P_{2}^{i\nu}(u_{0})\right)R^{2}\>,\,C=-\frac{i}{4\alpha g\gamma R^{2}}\>,$ (4.61)
as does continuity,
$B=-\frac{i\pi}{2\gamma R^{2}}\operatorname{csch}(\nu\pi)\frac{W\left(P_{2}^{-i\nu}(u_{0}),F_{-i\alpha}(z_{0})\right)}{W\left(F_{-i\alpha}(z_{0}),P_{2}^{i\nu}(u_{0})\right)}\>,\,D=-\frac{i}{4\alpha g\gamma R^{2}}\frac{W\left(P_{2}^{i\nu}(u_{0}),F_{i\alpha}(z_{0})\right)}{W\left(F_{-i\alpha}(z_{0}),P_{2}^{i\nu}(u_{0})\right)}\>.$ (4.62)
Thus, the scattering contributions to the Green’s function are
$G_{l,+}^{>}(s,s^{\prime};E)=-\frac{i\pi}{2\gamma R^{2}}\operatorname{csch}(\nu\pi)P_{2}^{i\nu}(u)\left(P_{2}^{-i\nu}(u^{\prime})+\frac{W\left(P_{2}^{-i\nu},F_{-i\alpha}\right)}{W\left(F_{-i\alpha},P_{2}^{i\nu}\right)}P_{2}^{i\nu}(u^{\prime})\right)\>,$ (4.63)
$G_{l,\pm}^{>}(s,s^{\prime};E)=\frac{P_{2}^{i\nu}(u)F_{-i\alpha}(z^{\prime})}{W\left(F_{-i\alpha},P_{2}^{i\nu}\right)R^{2}}\>,$ (4.64)
and
$G_{l,-}^{>}(s,s^{\prime};E)=-\frac{i}{4\alpha g\gamma R^{2}}F_{-i\alpha}(z^{\prime})\left(F_{i\alpha}(z)+\frac{W\left(P_{2}^{i\nu},F_{i\alpha}\right)}{W\left(F_{-i\alpha},P_{2}^{i\nu}\right)}F_{-i\alpha}(z)\right)\>.$ (4.65)
All Energies
The similarities between the oscillatory and attenuated expressions strongly hint towards a universal expression that is valid for all energies. Naively, starting from $n$ and increasing $E$ until it is greater than the sum of the angular momentum and mass terms yields $i\nu$, but the validity of such a statement depends on the branch taken. The branch ambiguity is resolved in the same way we resolved the issue of boundary conditions; by the $i\epsilon$ prescription.
Consider the expressions
$z=\lim_{\epsilon\rightarrow 0}\sqrt{c^{2}-(x^{2}+i\epsilon)}$ (4.66)
and
$w=\sqrt{x^{2}-c^{2}}\>,$ (4.67)
$z=\lim_{\epsilon\rightarrow 0}\sqrt{|c^{2}-x^{2}-i\epsilon|}\exp\left(\frac{i}{2}\operatorname{arg}\left(c^{2}-x^{2}-i\epsilon\right)\right)$ (4.68)
and
$w=\sqrt{|x^{2}-c^{2}|}\exp\left(\frac{i}{2}\operatorname{arg}\left(x^{2}-c^{2}\right)\right)\>.$ (4.69)
In the case that $x=0$, we have
$=c\>,$ (4.70)
and
$=ce^{i\pi/2}\>.$ (4.71)
For the case in which $0<x<c$, equivalently $c^{2}-x^{2}>0$, we have
$z=\sqrt{c^{2}-x^{2}}\>,$ (4.72)
while
$w=\sqrt{c^{2}-x^{2}}e^{i\pi/2}\>.$ (4.73)
$z=\sqrt{x^{2}-c^{2}}e^{-i\pi/2}\>,$ (4.74)
$w=\sqrt{x^{2}-c^{2}}\>.$ (4.75)
@@
where the limit $\epsilon\rightarrow 0$ is implicit. Thus, without loss of generality, we adopt the bound regime expressions hereafter.
All that remains is the sum over $l$, a task which cannot be completed in closed form but for which we determine an analytical approximation. In the so-called planar-wall approximation [40], we consider the source to be so large that its surface can be taken to be a plane. Let $z_{\perp}$ be the coordinate perpendicular to the surface and $\mathbf{z}_{\parallel}$ be a vector containing the coordinates which lie within it. We may Fourier transform with respect to the parallel coordinates, introducing a 2-momentum $\mathbf{k}$ in the process,
@@
writing $z_{\perp}=z$ for clarity. The 2-momentum components $G(z,z^{\prime};\mathbf{k},E)$ satisfy the equation
@@
which is just the equation for $R^{2}G_{l}(s,s^{\prime};E)$ but with the discrete parameter $l(l+1)/R^{2}$ replaced by the continuous parameter $k^{2}$. We will adopt the notation $G(s,s^{\prime};E,k)$ to reflect this change. It will turn out that the differential operator which defines $\delta\phi$ picks out a particular value of $k^{2}$ and $E^{2}$, so integrating over all modes is unnecessary at this point. However, we will revisit this point when we compute the tadpole contribution in the following section.

### v2

@@ 3.2 Thin-wall approximation
We want an analytical understanding of the quantum corrections, but the above system of equations cannot be solved in closed form. One could obtain an approximate solution by linearising the equation of motion [42], but then the Green’s function would be independent of the field. Instead, we consider a source whose radius $R$ is large compared to the characteristic scale of variations of the classical field. That is, $\mu R\gg 1$. Then, $\phi$ remains small and grows slowly in the interval $0<r<R$, experiences a rapid jump around $r=R$ and then quickly flattens out to $v$ as $r\rightarrow\infty$. Thus, we may drop the damping term in eq. (3.10) and turn what was once a radial problem into a purely one-dimensional problem. It is perhaps more conceptually clear to think of this approximation as an asymptotic approximation in the limit $\mu R\gg 1$. For notational clarity, we write s=rRs=r-R so that the equation of motion reads
@@
We denote by $\phi_{-}(s)$ and $\phi_{+}(s)$ the interior and exterior fields respectively. Separately solving for them is simply a matter of picking appropriate values for $s_{1}$ and $s_{2}$. The full profile
$\phi(s)=\Theta(-s)\phi_{-}(s)+\Theta(s)\phi_{+}(s)$ (3.17)
@@
for some real $c_{+}$.
For the interior contribution, choosing $s_{1}=$-$\infty$ and $s_{2}=s<0$ implies
@@
$g=\sqrt{2\left(\frac{\rho_{0}}{\mu^{2}M^{2}}-1\right)}$ (3.25)
is the effective mass of interior scalar fluctuations in units of $\gamma$. By the identity
@@
with $c_{-}\in\mathbb{R}$.
Continuity with $\chi_{+}$ is easily achieved by defining the constants $c_{\pm}$ in terms of the value of the field at the boundary of the source $\chi_{0}$, i.e.
@@
In conclusion, the static classical field profile is
$\varphi_{\text{cl}}(s)=v\left\{\Theta(-s)g\,\text{csch}\>\!\left[\text{arcsch}\>\!\left(\frac{\chi_{0}}{g}\right)-g\gamma s\right]+\Theta(s)\tanh\left(\gamma s+\text{artanh}\>\!\chi_{0}\right)\right\}$ (3.30)
@@ 4 Green’s function
We now turn to the calculation of the Green’s function. By means of the thin-wall approximation, we again reduce the spherical problem to a one-dimensional one. As one would expect from the problem of eigenmodes of a system with a piecewise continuous mass (see, for example, appendix B.1), the linearly independent solutions which compose the Green’s function behave differently for different frequency intervals. For frequencies below the mass of exterior scalar fluctuations, the solutions asymptotically vanish. Between the masses of interior and exterior fluctuations, the interior modes vanish asymptotically whereas the exterior modes oscillate. Above the interior mass, both interior and exterior modes oscillate unattenuated. However, the expressions for each frequency can be rendered equivalent by a Feynman contour prescription (see appendix B.2 for details).
The differential operator from which the Green’s function derives is related to the second functional derivative evaluated on $\varphi_{\text{cl}}$, slightly altered from the more general eq. (2.10) to include the effective potential,
@@
We again refer to the operator in parentheses as the fluctuation operator,
@@
where $\theta$ is the angle between $\mathbf{x}$ and $\mathbf{x}^{\prime}$ and $P_{l}$ is the Legendre polynomial of order $l$, allowing us to write
@@
Substituting this into eq. (4.7) and taking the thin-wall approximation (which amounts to dropping the damping term and replacing $r\rightarrow R$ in the centrifugal potential and discontinuity [36]) yields
@@
Details for how we solve for $G_{l}(s,s^{\prime};E)$ are given in appendix B.2. We summarise the results here. Let $u(s)=\tanh(\gamma s+c_{+})$ and $z(s)=\text{csch}\>\!^{2}(c_{-}-g\gamma s)$. Furthermore, define the parameters
@@
with $\epsilon>0$, where the $i\epsilon$ prescription ensures the following expressions are valid for all $E$. When both $s$ and $s^{\prime}$ are positive, the solution is
$G_{l,+}^{>}(s,s^{\prime};E)=-\frac{\pi}{2\gamma R^{2}}\csc\left(n\pi\right)P_{2}^{-n}(u)\left(P_{2}^{n}(u^{\prime})+\frac{W\left(P_{2}^{n}(u),F_{a}(z)\right)}{W\left(F_{a}(z),P_{2}^{-n}(u)\right)}P_{2}^{-n}(u^{\prime})\right)\>,$ (4.13)
where $W$ denotes the Wrońskian with respect to $s$, evaluated at $s=0$. $P_{2}^{n}$ is the associated Legendre function of the first kind and $F_{a}$ is related to the hypergeometric function $_{2}F_{1}$ by
$F_{a}(z)=z^{a}{}_{2}F_{1}\left(a-1,a+\frac{3}{2};2a+1;-z\right)\>.$ (4.14)
When both $s$ and $s^{\prime}$ are negative, we denote the solution by
$G_{l,-}^{>}(s,s^{\prime};E)=-\frac{1}{4ag\gamma R^{2}}F_{a}(z^{\prime})\left(F_{-a}(z)+\frac{W\left(P_{2}^{-n}(u),F_{-a}(z)\right)}{W\left(F_{a}(z),P_{2}^{-n}(u)\right)}F_{a}(z)\right)\>.$ (4.15)
Finally, when the signs of $s$ and $s^{\prime}$ differ, the Green’s function $G_{l}(s,s^{\prime};E)$ is given by
$G^{>}_{l,\pm}(s,s^{\prime};E)=\frac{P_{2}^{-n}(u)F_{a}(z^{\prime})}{W(F_{a}(z),P_{2}^{-n}(u))R^{2}}\>.$ (4.16)
The superscript >> corresponds to $s>s^{\prime}$. Solutions for $s<s^{\prime}$ are obtained from the relation $G_{l}^{<}(s,s^{\prime};E)=G_{l}^{>}(s^{\prime},s;E)$.
All that remains is the sum over $l$, a task which cannot be completed in closed form but for which we determine an analytical approximation. In the so-called planar-wall approximation [36], which is compatible with the thin-wall approximation, we consider the source to be so large that its surface can be taken to be a plane. Let $z_{\perp}$ be the coordinate perpendicular to the surface and $\mathbf{z}_{\parallel}$ be a vector containing the coordinates which lie within it. We may Fourier transform with respect to the parallel coordinates, introducing a 2-momentum $\mathbf{k}$ in the process,
@@
writing $z_{\perp}=z$ for ease. The function $G(z,z^{\prime};\mathbf{k},E)$ satisfies the equation
@@
which is just the equation for $R^{2}G_{l}(s,s^{\prime};E)$ but with the discrete parameter $l(l+1)/R^{2}$ replaced by the continuous parameter $k^{2}$. We will adopt the notation $G(s,s^{\prime};E,k)$ to reflect this change.